OpenStudy (anonymous):

Given the equation Square root of 2x plus 1 = 3, solve for x and identify if it is an extraneous solution. x = 4, solution is extraneous x = 4, solution is not extraneous x = 5, solution is extraneous x = 5, solution is not extraneous

OpenStudy (anonymous):

It's 4 but how do you know if it's extraneous or not?

OpenStudy (anonymous):

Plug it back into the original equation. If you get \(\sqrt{\text{negative number}}\), it's an extraneous solution.

OpenStudy (anonymous):

2(4) + 1 = 3? 2(4) + 1 = 9? So it would be?

OpenStudy (anonymous):

Think I'm doin something wrong :p

OpenStudy (anonymous):

No, it's not extraneous: \[\sqrt{2x+1}~~\Rightarrow~~\sqrt{8+1}=\sqrt{9}=\sqrt{\text{positive number}}\]

OpenStudy (anonymous):

Ok I get it, thanks

OpenStudy (anonymous):

yw